# Yang-Lee Quantum Criticality in Various Dimensions

### v1

@@ 5.1 Fuzzy sphere approach to Yang-Lee quantum criticality
In this section we use the fuzzy sphere regulator for $S^{2}$. This is a novel setup using the Quantum Hall system. It is analogous to Quantum Spin systems but preserves the SO(3) symmetry. The model describes non-relativistic fermions projected to the lowest Landau Level (LLL) with quadratic coupling and quartic couplings[13]
@@
where $\mathbf{c}_{m_{1}}=\begin{pmatrix}c_{m,\uparrow}\\ c_{m,\downarrow}\end{pmatrix}$. The the coefficient tensor
@@
is parametrized by Haldane pseudo-potential $V_{\ell}$. In our work we truncate the potential to $V_{0}$ and $V_{1}$, which is equivalent to the derivative expansion truncated to order $\nabla^{2}$.The overall normalization is fixed with $V_{1}=1$ and $V_{0}$ is a tunable parameter of the model. The LLL-projected Hamiltonian has $(4s+2)$ levels. We fill a the total number of $N=2s+1$ fermions in these levels. Under half-filling the charge degrees of freedom are gapped, and also the particle-hole symmetry is preserved. The total number of spins $N$ can be treated as a UV cutoff, which can be most clearly seen from the fact that the commutator $\{\psi(\Omega)\psi(\Omega^{\prime})\}$ approaches a delta function of size $\sqrt{1/N}$. This allows us to identify the radius of the sphere as $R\sim\sqrt{N}$.
@@ 5.2 Approaching the 3D YL critical point
The identification of each state is taken from the operator flow from Ising to Yang-Lee, shown in Fig. 8 and Table 9. Notice, that the expression (5.7) does not depend on the scaling dimension of $\phi$, which we will later find. From this, we obtain $ih_{z}^{*}(N)$ for $N=8,9,\dots,14$ and the IR critical point is extrapolated via (3.3) using a $1/N$ fitting. In Fig. 9, we show an example of the thermodynamic critical point extrapolation as a function on $N$ for distinct criticality criteria [^4] For the FSS criterion, we followed [58] and defined $h_{z}^{*}(N)$ as the point where the function $\sqrt{N+1}E_{10}^{(\textbf{1})}(N+1)=\sqrt{N}E_{10}^{(\textbf{1})}(N),$ here we have used that $R\propto\sqrt{N}$ [64]. , we can see they all converge to the same value within $\sim 1\%$, and we notice that $ih_{z}^{\textrm{crit}}$ is far from its pseudo-critical couterpart, in contrast to the 2D case, where $ih_{z}^{\textrm{crit}}$ and $ih_{z}^{*}(N)$ lie within a small perturbative regime [61]. We present some of the extrapolated critical points for various $h_{x}$ in Table 10 and we observe that the deeper we go into paramagnetic phase ($h_{x}>3.16$), the better the convergence; however, if we take a very big $h_{x}$ then we start facing numerical convergence problems. Therefore, we chose $h_{x}\in[4,100]$.
@@ 5.3 Extracting CFT data
CFT systems have many symmetries. In the fuzzy sphere realization this means each conformally invariant quantity has infinite many ways to extract. We see remarkable agreement between many distinct ways to evaluate the same quantity, which provide us non-trivial validations of our numeric procedure. In the bulk of this section we will walk through them in details.
Observable Fuzzy sphere 5-loop Two-sided 5-loop Truncated High-temperature
Padé Padé All [30] Bootstrap[32] expansion[41]
@@
$\Delta_{Q^{\mu\nu\kappa\lambda}}$ $4.9(1)_{[E]}$ $4.519_{[1,1]}$ $4.681_{[1,2]}$ $4.75(1)$
@@ 5.3.1 Scaling dimension
In radial quantization, the CFT states and local operators has one-to-one correspondence. The CFT Hamiltonian is the dilatation operator, so CFT local operators correspond to eigenstates and the associated eigenvalues are the scaling dimension. Each eigenstate also diagonalize global symmetry quantum numbers, which are total spin $\ell$, spin-$z$ projection $\ell_{z}$, and the spacetime parity $p$. The ground state has always spin 0.
In the fuzzy sphere realization, the energy levels can be shifted by an arbitrary ground state energy and rescaled by a model-dependent speed of light. Therefore, we will read off the CFT scaling dimensions as ratios. We will assume the lowest spin-0 state, the 2nd lowest spin-0 state, and the lowest spin-1 state correspond to identity $\mathbbm{I}$, the fundamental primary operator $\phi$, and its descendant $\partial\phi$ operators, respectively. The difference between the scaling dimensions of the primary and its first descendant is $\Delta_{\partial\phi}-\Delta_{\phi}=1$, which provides us a simple way to rescale the energy eigenvalues to match the CFT scaling dimensions, as in (5.7). For each state, the scaling dimension of the corresponding local operator is
@@
where ${\cal O}_{I}$ represents the leading irrelevant deformation. This means the contribution from the relevant deformation and the leading irrelevant deformation are at the same order of magnitude, assuming $g_{I}$, the dimensionless coupling constant of the irrelevant deformation is order 1. From this starting point, we see that the eigenvalues determined using the ratios (5.9) will contain finite size error as
@@
where $\Delta$ is the scaling dimension prediction extrapolated to the thermodynamic limit. The model is essentially a polynomial in $\frac{1}{R}=\frac{1}{\sqrt{N}}$ except for beginning at power $w$. We choose this model because it take into account that the leading correction comes in a specific power $N^{(3-\Delta_{{\cal O}_{I}})/2}$ in conformal perturbation theory, and the series expansion after the leading term is agnostic of the precise form of subleading irrelevant deformations and higher order perturbation theory, which we do not know precisely. For the 3D Yang-Lee criticality, it is likely that the irrelevant deformation ${\cal O}_{I}$ is the subleading scalar primary operator, which has a Ginzburg-Landau description $\phi^{3}$. From $(6-\epsilon)$ expansion and from our fuzzy sphere data we see that $\Delta_{\phi^{3}}\approx 4.6$, and we set $w=0.8$. The extrapolations from the fit should not be, and are not sensitive to the precise value of the input power $w$. For the number of powers we keep, we use $K=3$ which works most energy levels except for a few. We detect that $K$ is too high if the coefficient of the powers abruptly increases. To validate the fitting scheme, we test it on the difference between two levels $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ which we expect to be 2. A linear extrapolation is still visibly deviating from $2$. A higer power series fit brings us significantly closer to the theoretical value. The precise value of the power does not matter very much, so we extrapolation from various powers $w$ and truncations $K$ agree to the third digit, though $w=0.8$, the more informed guess from the conformal perturbation theory, is more stable against small change in truncation power.
In Fig. 11 we show the low spectrum of 3D YL criticality as a result of the extrapolation. In the extrapolation scheme (5.14) we treat each eigenvalues independently, so the conformal algebra is not taken as an input except for rescaling the gap between $\Delta_{\phi}$ and $\Delta_{\partial\phi}$ to 1. In the plot we see that relation between $\Delta_{\phi}$ and the descendant levels are precisely integers, making it easy to distinguish primary levels from descendant levels. The scaling dimension of the stress tensor is fixed at $\Delta_{T_{\mu\nu}}=3$ by the criticality criterion, and all its descendants in this plot are integers to high precision. The plot shows two other primaries, one at $(\ell=0,\Delta\approx 4.6)$ and the other at $(\ell=4,\Delta\approx 4.9)$. The former agrees with the ($6-\epsilon$) expansion prediction of the $\phi^{3}$ operator in the Ginzburg-Landau description, and the latter seems to belong to the leading double trace family $[\phi\phi]_{0,\ell}$ at spin 4, where the stress tensor also reside. We will denote the spin 4 primary as $Q_{\mu\nu\kappa\lambda}$.
@@
It is worth emphasizing that our confidence in the fuzzy sphere realization of the quantum criticality and the thermodynamic limit extrapolation is increased by the high precision agreement with conformal symmetry, which is presented in Fig. 13. The fact that the descendants of different spins agree with each other and are separated by integers is a result of the conformal algebra. We see that the agreement generically is broken for all states by a significant amount at finite size, but the extrapolation brings them to the corresponding integers in a non-linear manner. The agreement across many states is reached without any fine-tuning of fitting parameters. In Fig. 14, we utilize this property to determine the scaling dimension of the spin-4 primary $Q^{\mu\nu\kappa\lambda}$. As we go to the large spin regime we inevitably find a denser spectrum. Indeed, there are 3 operators within a small domain $4.214\leq\Delta\leq 5$: $\partial^{\mu}\partial^{\nu}\partial^{\kappa}\partial^{\lambda}\phi$, $\partial^{\mu}\partial^{\nu}T^{\kappa\lambda}$, and $Q^{\mu\nu\kappa\lambda}$. As a result, the fit to $\Delta_{Q^{\mu\nu\kappa\lambda}}$ is unstable. Instead, the spacetime parity-odd sector has a cleaner spectrum: there is no anti-symmetric descendants of $\phi$, and the leading operator in the spin-4 parity odd sector is the anti-symmetric descendant of $Q$, $\partial^{[\mu}Q^{\sigma]\nu\kappa\lambda}$. This operator has a healthier extrapolation giving $\Delta_{\partial Q}=5.9(1)$, and by conformal symmetry it is expected to be exactly 1 above the primary scaling dimension. We conclude that
@@
Figure 15: Scaling dimension $\Delta_{\phi}$ (left panel) and $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ (right panel) at the critical point of systems with various $h_{x}$ coupling. The prediction from $N\rightarrow\infty$ extrapolation is stable for a wide range of $h_{x}$, indicating that they belong to the same universality class. The $\Delta_{\phi}$ prediction from $h_{x}=4$ deviates from the rest of the result by $\sim 0.05$, indicating that small $h_{x}$ may be converging much slower than large $h_{x}$.
Yang-Lee criticality has 1 relevant parameter, which means in a phase diagram of more than 1 parameters it will appear as a the critical line or plane. In the phase diagram proposed in Fig. 4, we obtain a critical line where there is a critical coupling $ih_{z,c}$ for each $h_{x}$. This means we should obtain the same criticality for different $h_{x}$ and the extracted critical data should agree. In Fig. 15 we test the agreement of $\Delta_{\phi}$ and conformal algebra prediction $\Delta_{\partial^{2}\phi}-\Delta_{\phi}=2$ for a range of $h_{x}$. The predictions from different $h_{x}$ all agree when extrapolated to the thermodynamic limit, realizing the critical universality. We also see that large $h_{x}$ is computationally favorable as it seem to receive smaller finite size error.
@@ 5.3.2 OPE coefficients
$M_{\ell}^{a}$ $=\frac{1}{2s+1}\int d^{2}\Omega Y_{\ell 0}\psi^{\dagger}(\Omega)\sigma^{a}\psi(\Omega),\quad X_{\ell}\equiv n_{\ell}^{x},\quad Z_{\ell}\equiv n_{\ell}^{z}$ (5.18)
@@
$\langle{\cal O}(x_{1}){\cal O}(x_{2})\rangle$ $=\frac{1}{x_{12}^{2\Delta_{1}}}$ (5.20)
@@
$\langle 0|M_{0}^{a}|0\rangle=b_{0}\delta_{{\cal O}_{1}{\cal O}_{3}}+b_{1}C_{{\cal O}_{1}\phi{\cal O}_{3}}+\cdots$ (5.23)
@@
$\Delta_{\phi}=6\frac{\langle\partial\phi|M_{0}^{a}|\partial\phi\rangle-\langle 0|M_{0}^{a}|0\rangle}{\langle\phi|M_{0}^{a}|\phi\rangle}-3$ (5.30)
@@
It is remarkable that the $\Delta_{\phi}$ determined this way has significantly smaller finite size error. Indeed, the subleading terms in the $M_{0}^{a}$ has a special feature. We expand it more to include the descendants of $\phi$
@@
$M_{0}^{a}=b_{a,0}\mathbbm{1}+\bigg{(}b_{a,1}+(-1)^{k}\sum_{k=2}b_{a,1}^{(k)}(\Delta_{\phi})_{k}\bigg{)}\int d^{2}\Omega\,Y_{00}(\Omega)\phi(\Omega)+{(\text{more primaries}\cdots)}~{}.$ (5.34)
@@
$V_{i}$ $=\frac{x_{ij}^{2}(x_{ik}\cdot z_{i})-x_{ik}^{2}(x_{ij}\cdot z_{i})}{x_{jk}^{2}}$
In this basis, the correspondence between the fuzzy sphere matrix elements and the corresponding OPE coefficient is
$\langle{\cal O},\ell|n_{s}^{a}|{\cal O}^{\prime},\ell^{\prime}\rangle=\delta_{{\cal O},{\cal O}^{\prime}}\delta_{s,0}b_{0}+b_{a,1}f_{\ell,\ell^{\prime},s}C_{{\cal O}\phi{\cal O}^{\prime}}^{(s)}+\cdots,$ (5.38)
where $f_{\ell,\ell^{\prime},s}$ is a exactly computable prefactor that only depends on the spins. The detailed derivation of this correspondence is included in Appendix B. With this convention we study the $\langle\phi TT\rangle$ correlator, which has three polarizations $s=0,1,2$. The result is shown in Fig. 18. For $s=0$, the convergence is good, and by extrapolating to the thermodynamics we obtain
@@
The error bar of the quantities is determined by the difference between the linear and quadratic fit. The fact that the two results differ by a gap greater than the error bar means the higher powers in $\frac{1}{N}$ still contributes significantly. As $s$ become larger, the convergence becomes worse. A remarkable feature of the stress tensor correlator is that the stress tensor is a conserved current, $\partial_{\mu}T^{\mu\nu}=0$, which generates additional equations of motion that relates the OPE coefficients of different polarizations. It turns out that this correlator only has one independent OPE coefficient[71]
@@
Figure 19: Picture of 24-cell as graph on the $F_{4}$ Coxeter plane.
@@
where the first sum is taken over all the $96$ edges and the last two sums go over the $24$ vertices.
The symmetry group of $24$-cell is the Weyl group of $F_{4}$, denoted as $W(F_{4})$. This group has $1152$ elements, $25$ conjugacy classes and $25$ irreps [72, 73, 74]:
@@
where the number represents an irrep’s dimension. For convenience we present the character table of $W(F_{4})$ in the Appendix A. The $O(4)$ irreps $(0,0)$, $(1/2,1/2)$ and $(1,1)$ do not split on 24-cell and correspond to irreps 1, 4 and 9. The $O(4)$ irrep $(3/2,3/2)$ splits into $\textbf{8}+\textbf{8}^{\prime}$ and the irrep $(2,2)$ splits into $\textbf{2}+\textbf{2}^{\prime}+\textbf{9}+\textbf{12}$.
We plotted the lower energy gaps of the Hamiltonian (6.1) along the path $(h_{x},ih_{z})=(0,0)\to(15,0)\to(5,i6)$ with $J=1$ in the Figure 20, where different colors correspond to different irreps of $W(F_{4})$.
Figure 20: Energy gaps of the Hamiltonian (6.1) for 24-cell along the path $(h_{x},ih_{z})=(0,0)\to(15,0)\to(5,i6)$, with $J=1$. We identified the lower energy levels with their corresponding operators in the Ising (Gaussian) and Yang-Lee CFTs. The dashed line is the pseudo-critical point $h_{z}^{*}$ obtained fixing $r_{T}=4$.
The behavior of the energy levels is similar to the 2D and 3D cases, and we can easily identify a few low-lying operators. We note that the 4D Ising CFT is a Gaussian fixed point in $d=4$; for example, the operators $\phi$, $\phi^{2}$ should have dimensions $\Delta_{\phi}=1$, $\Delta_{\phi^{2}}=2$. However, due to finite-size corrections for the Ising model on the 24-cell, the energy gaps are shifted and we do not observe Gaussian fixed-point behavior for most of the levels, nevertheless one can notice that there is a pseudo-critical point $h_{x}^{*}\approx 7.5$ where $\Delta_{\phi^{2}}\approx\Delta_{\partial\phi}\approx 2\Delta_{\phi}$. It would be interesting to apply Conformal Perturbation Theory to this model to better observe the Gaussian CFT behavior.
@@
There are five other regular 4D polytopes one can consider: the 5-cell, 16-cell, 8-cell, 120-cell, and 600-cell. The 8-cell (Hypercube) has only $N=16$ vertices, so the 4D Yang–Lee or 4D Ising (Gaussian) models can be easily analyzed numerically on this polytope via exact diagonalization. On the other hand, the 120-cell and 600-cell have $120$ and $600$ vertices, respectively, making exact diagonalization infeasible. However, modern numerical methods such as Neural Quantum States [75] could potentially be used to analyze the 4D Yang–Lee or 4D Ising (Gaussian) models on these large regular 4D polytopes. We leave this for future work.

### v2

@@ 5.1 Fuzzy sphere approach to Yang-Lee quantum criticality
In this section we use the fuzzy sphere regulator for $S^{2}$. This is a novel setup using the Quantum Hall system. It is analogous to Quantum Spin systems but preserves the SO(3) symmetry. The model describes non-relativistic fermions projected to the lowest Landau Level (LLL) with quadratic and quartic couplings[14, 22]
@@
where $\mathbf{c}_{m_{1}}=\begin{pmatrix}c_{m,\uparrow}\\ c_{m,\downarrow}\end{pmatrix}$. The coefficient tensor
@@
is parametrized by Haldane pseudo-potential $V_{\ell}$. In our work we truncate the potential to $V_{0}$ and $V_{1}$, which is equivalent to the derivative expansion truncated to order $\nabla^{2}$. The overall normalization is fixed with $V_{1}=1$ and $V_{0}$ is a tunable parameter of the model. The LLL-projected Hamiltonian has $(4s+2)$ levels. We fill a the total number of $N=2s+1$ fermions in these levels. Under half-filling the charge degrees of freedom are gapped, and also the particle-hole symmetry is preserved. The total number of spins $N$ can be treated as a UV cutoff, which can be most clearly seen from the fact that the commutator $\{\psi(\Omega),\psi(\Omega^{\prime})\}$ approaches a delta function of size $\sqrt{1/N}$. This allows us to identify the radius of the sphere as $R\sim\sqrt{N}$.
@@ 5.2 Approaching the 3D YL critical point
The identification of each state is taken from the operator flow from Ising to Yang-Lee, shown in Fig. 8 and Table 9. Notice, that the expression (5.7) does not depend on the scaling dimension of $\phi$, which we will later find. From this, we obtain $ih_{z}^{*}(N)$ for $N=8,9,\dots,14$ and the IR critical point is extrapolated via (3.3) using a $1/N$ fitting. In Fig. 9, we show an example of the thermodynamic critical point extrapolation as a function of $N$ for distinct criticality criteria [^4] For the FSS criterion, we followed [71] and defined $h_{z}^{*}(N)$ as the point where the function $\sqrt{N+1}E_{10}^{(\textbf{1})}(N+1)=\sqrt{N}E_{10}^{(\textbf{1})}(N),$ here we have used that $R\propto\sqrt{N}$ [76]. , we can see they all converge to the same value within $\sim 1\%$, and we notice that $ih_{z}^{\textrm{crit}}$ is far from its pseudo-critical couterpart for small $N$, in contrast to the 2D case, where $ih_{z}^{\textrm{crit}}$ and $ih_{z}^{*}(N)$ lie within a small perturbative regime [74]. We present some of the extrapolated critical points for various $h_{x}$ in Table 10 and we observe that the deeper we go into paramagnetic phase ($h_{x}>3.16$), the better the convergence; however, if we take a very big $h_{x}$ then we start facing numerical convergence problems. Therefore, we chose $h_{x}\in[4,100]$.
@@ 5.3 Extracting the CFT data
CFT systems have many symmetries. In the fuzzy sphere realization this means each conformally invariant quantity has infinitely many ways to be extract. We see remarkable agreement between many distinct ways to evaluate the same quantity, which provide us non-trivial validations of our numeric procedure. In the bulk of this section we will walk through them in details.
Observable Fuzzy sphere Padé Two-sided 5-loop Truncated High-temperature
Padé All [41] Bootstrap[43] expansion[52]
@@
$\Delta_{Q_{\mu\nu\kappa\lambda}}$ $4.9(1)_{[E]}$ $4.519_{[1,1]}$ $4.681_{[1,2]}$ $4.75(1)$
@@ 5.3.1 Scaling dimension
In radial quantization, the CFT states and local operators have one-to-one correspondence. The CFT Hamiltonian is the dilatation operator, so CFT local operators correspond to eigenstates and the associated eigenvalues are the scaling dimension. Each eigenstate also diagonalize global symmetry quantum numbers, which are total spin $\ell$, spin-$z$ projection $\ell_{z}$, and the spacetime parity $p$. The ground state has always spin 0.
In the fuzzy sphere realization, the energy levels can be shifted by an arbitrary ground state energy and rescaled by a model-dependent speed of light. Therefore, we will read off the CFT scaling dimensions as ratios. We will assume the lowest spin-0 state, the 2nd lowest spin-0 state, and the lowest spin-1 state correspond to identity I, the fundamental primary operator $\phi$, and its descendant $\partial\phi$ operators, respectively. The difference between the scaling dimensions of the primary and its first descendant is $\Delta_{\partial\phi}-\Delta_{\phi}=1$, which provides us a simple way to rescale the energy eigenvalues to match the CFT scaling dimensions, as in (5.7). For each state, the scaling dimension of the corresponding local operator is
@@
where ${\cal O}_{I}$ represents the leading irrelevant deformation. This means the contribution from the relevant deformation and the leading irrelevant deformation are at the same order of magnitude, assuming the dimensionless coupling constant of the irrelevant deformation $g_{I}$ is of order 1. From this starting point, we see that the eigenvalues determined using the ratios (5.9) will contain finite size error as
@@
where $\Delta$ is the scaling dimension prediction extrapolated to the thermodynamic limit. The model is essentially a polynomial in $\frac{1}{R}=\frac{1}{\sqrt{N}}$ except for beginning at power $w$. We choose this model because it takes into account that the leading correction comes in a specific power $N^{(3-\Delta_{{\cal O}_{I}})/2}$ in conformal perturbation theory, and the series expansion after the leading term is agnostic of the precise form of subleading irrelevant deformations and higher order perturbation theory, which we do not know precisely. For the 3D Yang-Lee criticality, it is likely that the irrelevant deformation ${\cal O}_{I}$ is the subleading scalar primary operator, which has a Ginzburg-Landau description $\phi^{3}$. From $(6-\epsilon)$ expansion and from our fuzzy sphere data we see that $\Delta_{\phi^{3}}\approx 4.6$, and we set $w=0.8$. The extrapolations from the fit should not be, and are not, sensitive to the precise value of the input power $w$. For the number of powers we keep, we use $K=3$ which works for most energy levels except for a few. We detect that $K$ is too high if the coefficient of the powers abruptly increases. To validate the fitting scheme, we test it on the difference between two levels $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ which we expect to be 2. A linear extrapolation is still visibly deviating from $2$. A higer power series fit brings us significantly closer to the theoretical value. The precise value of the power does not matter very much, so we extrapolation from various powers $w$ and truncations $K$ agree to the third digit, though $w=0.8$, the more informed guess from the conformal perturbation theory, is more stable against small change in truncation power.
In Fig. 11 we show the low spectrum of 3D YL criticality as a result of the extrapolation. In the extrapolation scheme (5.14) we treat each eigenvalues independently, so the conformal algebra is not taken as an input except for rescaling the gap between $\Delta_{\phi}$ and $\Delta_{\partial\phi}$ to 1. In the plot we see that relation between $\Delta_{\phi}$ and the descendant levels are precisely integers, making it easy to distinguish primary from descendant levels. The scaling dimension of the stress tensor is fixed at $\Delta_{T_{\mu\nu}}=3$ by the criticality criterion, and all its descendants in this plot are integers to high precision. The plot shows two other primaries, one at $(\ell=0,\Delta\approx 4.6)$ and the other at $(\ell=4,\Delta\approx 4.9)$. The former agrees with the ($6-\epsilon$) expansion prediction of the $\phi^{3}$ operator in the Ginzburg-Landau description, and the latter seems to belong to the leading double trace family $[\phi\phi]_{0,\ell}$ at spin 4, where the stress tensor also reside. We will denote the spin 4 primary as $Q_{\mu\nu\kappa\lambda}$.
@@
It is worth emphasizing that our confidence in the fuzzy sphere realization of the quantum criticality and the thermodynamic limit extrapolation is increased by the high precision agreement with conformal symmetry, which is presented in Fig. 13. The fact that the descendants of different spins agree with each other and are separated by integers is a result of the conformal algebra. We see that the agreement generically is broken for all states by a significant amount at finite size, but the extrapolation brings them to the corresponding integers in a non-linear manner. The agreement across many states is reached without any fine-tuning of fitting parameters. In Fig. 14, we utilize this property to determine the scaling dimension of the spin-4 primary $Q^{\mu\nu\kappa\lambda}$. As we go to the large spin regime we inevitably find a denser spectrum. Indeed, there are 3 operators within a small domain $4.214\leq\Delta\leq 5$: $\partial^{\mu}\partial^{\nu}\partial^{\kappa}\partial^{\lambda}\phi$, $\partial^{\mu}\partial^{\nu}T^{\kappa\lambda}$, and $Q^{\mu\nu\kappa\lambda}$. As a result, the fit to $\Delta_{Q^{\mu\nu\kappa\lambda}}$ is unstable. Instead, the spacetime parity-odd sector has a cleaner spectrum: there are not anti-symmetric descendants of $\phi$, and the leading operator in the spin-4 parity odd sector is the anti-symmetric descendant of $Q$, $\partial^{[\mu}Q^{\sigma]\nu\kappa\lambda}$. This operator has a healthier extrapolation giving $\Delta_{\partial Q}=5.9(1)$, and by conformal symmetry it is expected to be exactly 1 above the primary scaling dimension. We conclude that
@@
Yang-Lee criticality has 1 relevant parameter, which means in a phase diagram of more than 1 parameters it will appear as a critical line or plane. In the phase diagram proposed in Fig. 4, we obtain a critical line where there is a critical coupling $ih_{z,c}$ for each $h_{x}$. This means we should obtain the same criticality for different $h_{x}$ and the extracted critical data should agree. In Fig. 15 we test the agreement of $\Delta_{\phi}$ and conformal algebra prediction $\Delta_{\partial^{2}\phi}-\Delta_{\phi}=2$ for a range of $h_{x}$. The predictions from different $h_{x}$ all agree when extrapolated to the thermodynamic limit, realizing the critical universality. We also see that large $h_{x}$ is computationally favorable as it seem to receive smaller finite size error.
Figure 15: Scaling dimension $\Delta_{\phi}$ (left panel) and $\Delta_{\partial^{2}\phi}-\Delta_{\phi}$ (right panel) at the critical point of systems with various $h_{x}$ couplings. The prediction from $N\rightarrow\infty$ extrapolation is stable for a wide range of $h_{x}$, indicating that they belong to the same universality class. The $\Delta_{\phi}$ prediction from $h_{x}=4$ deviates from the rest of the result by $\sim 0.05$, indicating that small $h_{x}$ may be converging much slower than large $h_{x}$.
@@ 5.3.2 OPE coefficients
$M_{\ell}^{a}$ $=\frac{1}{2s+1}\int d^{2}\Omega Y_{\ell 0}\psi^{\dagger}(\Omega)\sigma^{a}\psi(\Omega),\quad X_{\ell}\equiv M_{\ell}^{x},\quad Z_{\ell}\equiv M_{\ell}^{z}$ (5.18)
@@
$\langle{\cal O}(x_{1}){\cal O}(x_{2})\rangle$ $=\frac{1}{x_{12}^{2\Delta}}$ (5.20)
@@
$\langle{\cal O}_{1},0|M_{0}^{a}|{\cal O}_{3},0\rangle=b_{0}\delta_{{\cal O}_{1}{\cal O}_{3}}+b_{1}C_{{\cal O}_{1}\phi{\cal O}_{3}}+\cdots$ (5.23)
@@
$\Delta_{\phi}=6\frac{\langle\partial\phi|M_{0}^{a}|\partial\phi\rangle-\langle 0|M_{0}^{a}|0\rangle}{\langle\phi|M_{0}^{a}|\phi\rangle-\langle 0|M_{0}^{a}|0\rangle}-3$ (5.30)
@@
It is remarkable that the $\Delta_{\phi}$ determined this way has significantly smaller finite size error. Indeed, the subleading terms in the $M_{0}^{a}$ have a special feature. We expand it more to include the descendants of $\phi$
@@
$M_{0}^{a}=b_{a,0}\mathbbm{1}+\bigg{(}b_{a,1}+\sum_{k=2}(-1)^{k}b_{a,1}^{(k)}(\Delta_{\phi})_{k}\bigg{)}\int d^{2}\Omega\,Y_{00}(\Omega)\phi(\Omega)+{(\text{more primaries}\cdots)}~{}.$ (5.34)
@@
$V_{i,j,k}$ $=\frac{x_{ij}^{2}(x_{ik}\cdot z_{i})-x_{ik}^{2}(x_{ij}\cdot z_{i})}{x_{jk}^{2}}~{}.$
In this basis, the correspondence between the fuzzy sphere matrix elements and the OPE coefficients is
$\langle{\cal O},\ell|M_{s}^{a}|{\cal O}^{\prime},\ell^{\prime}\rangle=\delta_{{\cal O},{\cal O}^{\prime}}\delta_{s,0}b_{0}+b_{a,1}f_{\ell,\ell^{\prime},s}C_{{\cal O}\phi{\cal O}^{\prime}}^{(s)}+\cdots,$ (5.38)
where $f_{\ell,\ell^{\prime},s}$ is an exactly computable prefactor that only depends on the spins. The detailed derivation of this correspondence is included in Appendix B. With this convention we study the $\langle\phi TT\rangle$ correlator, which has three polarizations $s=0,2,4$. The result is shown in Fig. 18. For $s=0$, the convergence is good, and by extrapolating to the thermodynamic limit we obtain
@@
The error bar of the quantities is determined by the difference between the linear and quadratic fit. The fact that the two results differ by a gap greater than the error bar means the higher powers in $\frac{1}{N}$ still contributes significantly. As $s$ become larger, the convergence becomes worse. A remarkable feature of the stress tensor correlator is that the stress tensor is a conserved current, $\partial_{\mu}T^{\mu\nu}=0$, which generates additional equations of motion that relate the OPE coefficients of different polarizations. It turns out that this correlator only has one independent OPE coefficient[83]
@@
Figure 19: Picture of the 24-cell as graph on the $F_{4}$ Coxeter plane.
@@
where the first sum is taken over all the $96$ edges and the last two sums go over the $24$ vertices. The symmetry group of $24$-cell is the Weyl group of $F_{4}$, denoted as $W(F_{4})$. This group has $1152$ elements, $25$ conjugacy classes and $25$ irreps [84, 85, 86]:
@@
where the number represents an irrep’s dimension. For convenience we present the character table of $W(F_{4})$ in the Appendix A. The $O(4)$ irreps $(0,0)$, $(1/2,1/2)$ and $(1,1)$ do not split on 24-cell and correspond to irreps 1, 4 and 9. The $O(4)$ irrep $(3/2,3/2)$ splits into $\textbf{8}+\textbf{8}^{\prime}$, $(2,2)$ splits into $\textbf{2}+\textbf{2}^{\prime}+\textbf{9}+\textbf{12}$ and $(5/2,5/2)$ splits into $\textbf{4}+\textbf{8}+\textbf{8}^{\prime}+\textbf{16}$.
We exhibit the lower energy gaps of the Hamiltonian (6.1) along the path $(h_{x},ih_{z})=(0,0)\to(15,0)\to(15,i6)$ with $J=1$ in the Figure 20, where different colors correspond to different irreps of $W(F_{4})$.
Figure 20: Energy gaps of the Hamiltonian (6.1) for 24-cell along the path $(h_{x},ih_{z})=(0,0)\to(15,0)\to(15,i6)$, with $J=1$. We identified the lower energy levels with their corresponding operators in the Ising (Gaussian) and Yang-Lee CFTs. The dashed line is the pseudo-critical point $h_{z}^{*}$ obtained fixing $r_{T}=4$.
The behavior of the energy levels is similar to the 2D and 3D cases, and we can easily identify a few low-lying operators. We note that the 4D Ising model flows to a Gaussian fixed point in $d=4$; for example, the operators $\phi$, $\phi^{2}$ should have dimensions $\Delta_{\phi}=1$, $\Delta_{\phi^{2}}=2$. However, due to finite-size corrections for the Ising model on the 24-cell, the energy gaps are shifted and we do not observe the Gaussian fixed-point behavior for most of the levels, nevertheless one can notice that there is a pseudo-critical point $h_{x}^{*}\approx 7.5$ where $\Delta_{\phi^{2}}\approx\Delta_{\partial\phi}\approx 2\Delta_{\phi}$. It would be interesting to apply Conformal Perturbation Theory to this model to better observe the Gaussian CFT behavior.
@@
There are five other regular 4D polytopes one can consider: the 5-cell, 16-cell, 8-cell, 120-cell, and 600-cell. The 8-cell (Hypercube) has only $N=16$ vertices, so the 4D Yang–Lee or 4D Ising (Gaussian) models can be easily analyzed numerically on this polytope via exact diagonalization. On the other hand, the 120-cell and 600-cell have $600$ and $120$ vertices, respectively, making exact diagonalization unfeasible. However, modern numerical methods such as Neural Quantum States [87] could potentially be used to analyze the 4D Yang–Lee or 4D Ising models (the latter flows to the Gaussian one) on these large regular 4D polytopes. We leave this for future work.